# Integrated population model reveals human and environment driven changes in Baltic ringed seal (Pusa hispida botnica) demography and behavior

### v1

@@ 4 Discussion
Our estimates suggest that the composition of the Swedish hunting bag may differ significantly between the spring and the fall. We found that hunting was largely aimed at adults during the fall, possibly because of the greater damage they cause to fishing gear and catch. Studies of grey seals have shown that some adults, typically males, may specialize in raiding fishing gear (konigson2013). In contrast, the adult bias was completely absent during the spring hunt. This is not particularly surprising considering that adult seals are molting on ice in large numbers during the spring, and seldom venture out at sea (smith1970; Harkonen1992; born2002; Kelly2010). Younger seals may be much more likely to be foraging in open water during most of the molting period, and thus have more frequent interactions with fishermen (smith1970; lydersen1993; crawford2012; Oksanen2015). In fact, we estimated that adult seals were generally more likely to be hunted than sub-adults during the spring in Finland, where hunting primarily takes place on sea ice. Pups were an exception and were overrepresented in the Finnish hunting bag despite their relatively low presence on sea ice, although some late born pups are often still with their mothers during the Finnish hunting season. Since hunting in Finland is recreational, hunters may specifically seek out inexperienced pups for their meat and fur (kingsley1998).
@@
We were not able to estimate the overall magnitude of bycatch mortality due to a lack of data. Such data could be obtained through interviews with fisherman (vanhatalo2014). However, consistent with findings from many other seal populations, we found that pups were significantly more at risk of getting bycaught in small fisheries than older seals (sipila1990; backlin2011; niemi2013; vanhatalo2014; jounela2019). As pups tend to favor shallow waters for foraging, and are generally curious and inexperienced, it is probable that they frequently engage with coastal fisheries and become caught in fishing gear (Kelly2010; niemi2013; cronin2014). Their relatively small body sizes may also make them more likely to enter smaller fishing gear, and make escaping more difficult once they are caught (cronin2014). The juxtaposition between the substantial juvenile bias in bycatch and the pronounced adult bias in protective hunting by fisherman may reflect the fact that young, inexperienced seals are inclined towards exploration, whereas experienced adults prioritize exploitation (sjoberg2000). On the other hand, larger seals may be more likely to detach from fishing nets, and fishermen often report the difficulty of lifting out larger seals from the water (cronin2014). Older seals may therefore be underrepresented in the samples compared to the actual bycatch.
@@ 5 Conclusions
Amid the dynamic and uncertain conditions brought on by climate change, varying reproductive rates, and the re-introduction of seal hunting, our Bayesian IPM succeeded in providing a comprehensive assessment of the Baltic ringed seal population in the Bothnian Bay—a task that has not been possible for over a decade. In addition to estimating population size and growth, we inferred key demographic parameters including survival probabilities, reproductive rates and the relative vulnerabilities of different demographic groups to hunting and bycatch. Moreover, our model revealed possible changes in ringed seal haul-out behavior in response to changing sea ice patterns. The mechanistic nature of our model additionally enabled near-term predictions of population response to changes in hunting quotas, informing management decisions.
@@ Acknowledgements
The present work was part of Murat Ersalman’s Master’s thesis. The work has received funding from the Helsinki Institute of Life Science (HiLIFE) (ME), Research Council of Finland (grant 317255) and Jane & Aatos Erkko Foundation (JV). The monitoring of ringed seals in the Bothnian Bay was conducted by the Swedish Museum of Natural History and funded by the Swedish Environmental Protection Agency and the Swedish Agency for Marine and Water Management. The Finnish monitoring of seals was funded by the Ministry of Agriculture and Forestry. We thank Yessenia Rojas, Jannikke Räikkönen, Petri Timonen and Charlotta Moraeus for help in age determinations, and Eva Kisdi for support in building the mathematical models.
@@
Figure 2: Illustration of the Baltic ringed seal life-cycle. The density plots at the center show, in polar coordinates, the temporal distribution of hunting and bycatch based on sampled seals. (The illustration was created with the aid of DALL-E 3.)
@@
Figure 4: Estimated population size LABEL:sub@fig:pop_size, birth rate LABEL:sub@fig:birth_rate, population growth rate LABEL:sub@fig:growth_rate and age structure LABEL:sub@fig:age_str for ringed seals in the Bothnian Bay. The rightmost segments show the expected asymptotic values for an unexploited population with and without density dependent effects.
Figure 5: LABEL:sub@fig:survival_probs Probabilities of surviving natural mortality (including bycatch), and the relative probabilities that seals are LABEL:sub@fig:bycatch_bias bycaught, LABEL:sub@fig:sw_bias_spring hunted in Sweden during the spring, LABEL:sub@fig:sw_bias_fall hunted in Sweden during the fall, and LABEL:sub@fig:fi_bias hunted in Finland. The dashed lines in LABEL:sub@fig:bycatch_bias-LABEL:sub@fig:fi_bias show the relative mortality probabilities if all individuals were equally likely to be hunted or bycaught. Calculation of the relative mortality probabilities are presented in Appendix S3.
@@
Figure 7: LABEL:sub@fig:critical_hunting Critical size of the hunting bag for which the population growth rate is expected to be zero, as a function of total population size. The critical size of the hunting bag is calculated such that the implied survival probabilities satisfy Equation (25) for $\lambda=1$. The red error bar shows the estimated population size and hunting quota (Finland and Sweden combined) in 2022. LABEL:sub@fig:future_preds Posterior predictions for future population growth rates under four different hunting quota scenarios: (1) quotas maintained at their current level of 350 and 375 seals in Sweden and Finland respectively, (2) quotas increased by a total of 30 licenses per year (15 in both Sweden and Finland), (3) quotas increased by a total of 70 licenses per year (35 in both Sweden and Finland) and (4) quotas reduced by a total of 540 in 2024 (270 in both Sweden and Finland) and kept constant thereafter.
Appendix Appendix S1 Posterior predictive checks
Figure S1: Posterior estimates and central 95% predictive intervals for the proportion of samples with a LABEL:sub@fig:preg_rate visible fetus, LABEL:sub@fig:scar_rate placental scar, and LABEL:sub@fig:CA_rate corpus albicans.
@@
Figure S2: Posterior predictive checks for the sizes of LABEL:sub@fig:hbag_sw Swedish and LABEL:sub@fig:hbag_fi Finnish hunting bags.
@@
Figure S3: Posterior predictive checks for samples from LABEL:sub@fig:h_samples_sw total Swedish hunting, LABEL:sub@fig:h_samples_fi Finnish hunting and LABEL:sub@fig:bycatch_samples bycatch. The density estimates are based on the frequency distribution of the number of sampled seals, pooled across all years and demographic classes.
Appendix Appendix S2 Prior distributions
@@
We parameterized the age and sex specific median hunting efforts, $\hat{E}_{s,a}$, as the product of a global scaling parameter, $\hat{E}$, and a vector of age-specific weights, $\boldsymbol{\psi}$, defined over the unit simplex, so that
@@
Decoupling the magnitude of hunting effort from its allocation to different demographic groups significantly improves sampling performance. Using Appendix S3: Equation S7, the scaling parameter $\hat{E}$ can be reparametrized as
@@
Annual variation in hunting effort was modeled as
@@
We chose a large standard deviation of $\sigma=5$ for all components of $\mathbf{g}$ (note that $\sigma$ is on a logarithmic scale in Equation S6), which yields a correlation coefficient of $\varrho=0.9887$.
Since the logistic transformation in Equation S6 requires only 11 degrees of freedom for the 12 demographic groups in our model, we arbitrarily set $g_{12}=0$, and computed the conditional covariance matrix:
@@
We can ensure that the maximum rate of change along the logistic curve in Equation (13) does not exceed the change in pregnancy rate implied by Equation S14 by satisfying the inequality
@@
Appendix S4: Equation S4 implies that $e^{-\theta_{0}}$ cannot be less than $b_{\text{max}}$. We therefore gave a Uniform($b_{\text{max}}$, 1) prior to $e^{-\theta_{0}}$. Instead of assigning a prior distribution to $\theta_{1}$, we parametrized our model in terms of the carrying capacity, $K$ (see Appendix S4: Equation S8). Estimates of the total number of Baltic ringed seals at the turn of ${}^{\text{th}}$ century range from 50,000 to 450,000 (durant1986; Harding1999; Kokko1999). However, estimates on the lower end of this range are based on very conservative assumptions, including the absence of density dependent regulation. The carrying capacity of the Baltic Sea is therefore highly likely to exceed 200,000 ringed seals (Harding1999). We assumed that, in the absence of hunting pressure, the carrying capacity of the Baltic Sea as a whole is likely to be between 200,000 and 450,000 ringed seals. Based on the distribution of sea ice across the Baltic Sea in late April, about 75% of the Baltic ringed seal population may be present in the Bothnian Bay during the aerial surveys. We therefore gave a Gamma(18, 8e-5) prior to the carrying capacity, $K$, of the Bothnian Bay, which has an expectation of 225,000 (and 95% CI: 133,000 - 340,000).
@@
Survival prob. of female pups $\phi_{f,0}$ $\frac{\phi_{f,0}}{\phi_{f,5+}}\sim U(0,1)$ 0.62 (0.41 - 0.87)
Deviation between mortality rates of male and female pups $\nu_{0}$ $\nu_{0}\sim N(0,0.1^{2})$ 0.01 (-0.18 - 0.21)
Deviation between mortality rates of 1+ yo male and females $\nu_{1+}$ $\nu_{1+}\sim N(0,0.1^{2})$ 0.00 (-0.20 - 0.19)
Shape param. for survival curve $c$ $c\sim U(0,1)$ 0.50 (0.03 - 0.97)
Scale of FI hunting effort $\hat{E}_{\text{fi}}$ $p_{Q}^{\text{fi}}\sim U(0,1)$ 0.45 (0.26 - 0.82)
Scale of spring SW hunting effort $\hat{E}_{\text{sw(1)}}$ $p_{Q}^{\text{sw(1)}}\sim U(0,1)$ 2e-3 (7e-4 - 4e-3)
Scale of fall SW hunting effort $\hat{E}_{\text{sw(2)}}$ $p_{Q}^{\text{sw(2)}}\sim U(0,1)$ 1e-3 (4e-4 - 4e-3)
SD for FI hunting effort $\sigma_{\text{E}_{\text{fi}}}$ $\sigma_{\text{E}_{\text{fi}}}\sim\text{Cauchy}_{+}(0,0.1)$ 0.48 (0.19 - 1.38)
SD for spring SW hunting effort $\sigma_{\text{E}_{\text{sw(1)}}}$ $\sigma_{\text{E}_{\text{sw(1)}}}\sim\text{Cauchy}_{+}(0,0.1)$ 1.0 (0.6 - 2.0)
SD for fall SW hunting effort $\sigma_{\text{E}_{\text{sw(2)}}}$ $\sigma_{\text{E}_{\text{sw(2)}}}\sim\text{Cauchy}_{+}(0,0.1)$ 1.2 (0.6 - 3.2)
@@
Ratio of pup to adult haul-out $\delta$ $\delta\sim U(0,1)$ 0.58 (0.19 - 0.95)
Intercept for haul-out model $\alpha_{0}$ $\frac{\alpha_{0}}{\alpha_{1}}\sim\text{Cauchy}(0,5)$ 17.6 (6.2 - 196.2)
Slope for haul-out model $\alpha_{1}$ $\alpha_{1}\sim\text{Cauchy}(0,1.25)$ 11.0 (4.4 - 121.9)
Half-saturation const. $f$ $f\sim\text{Cauchy}_{+}(0,37)$ 11.9 (0.9 - 23.3)
Ratio btw. asymptotes of movement rate into the water $d$ $d\sim U(0,1)$ 0.03 (0.00 - 0.20)
Max. birth rate $b_{\text{max}}$ $b_{\text{max}}\sim U(0,1)$ 0.79 (0.72 - 0.87)
Min. birth rate $b_{\text{min}}$ $\frac{b_{\text{min}}}{b_{\text{max}}}\sim U(0,1)$ 0.25 (0.02 - 0.52)
Intercept for logistic birth rate $\beta_{0}$ $\beta_{0}\sim\text{Cauchy}(0,10)$ 1.7 (-0.5 - 4.8)
Slope for logistic birth rate $\beta_{1}$ $\beta_{1}\sim\text{t}_{\nu=5}(0,1.8)$ 2.6 (1.0 - 7.1)
@@
Base pregnancy failure rate $\theta_{0}$ $e^{-\theta_{0}}\sim U(b_{\text{max}},1)$ 0.13 (0.01 - 0.26)
Dispersion parameter $r$ $\frac{1}{\sqrt{r}}\sim\text{Cauchy}_{+}(0,1)$ 32.9 (17.2 - 57.1)
Placental scar detection prob. $\pi_{\text{s}}$ $\pi_{\text{s}}\sim U(0,1)$ 0.92 (0.85 - 0.99)
CA detection prob. $\pi_{\text{c}}$ $\pi_{\text{c}}\sim U(0,1)$ 0.97 (0.91 - 0.99)
CA prob. for non-pregnant seals $\kappa$ $\kappa\sim U(0,1)$ 0.07 (0.00 - 0.24)
Sampling prob. for seals $<$100cm $x_{0}$ $x_{0}\sim U(0,1)$ See Appendix S5.
Sampling prob. for seals $>$100cm $x_{1}$ $x_{1}\sim U(0,1)$ See Appendix S5.
Probability that seal is $>$100cm $q_{s,a}$ $q_{s,a}\sim U(0,1)\ \ \forall s,a$ See Appendix S5.
Appendix Appendix S3 Hunting mortality

### v2

@@ 4 Discussion
Our estimates suggest that the demographic composition of Swedish harvests may differ significantly between the spring and the fall. We found that hunting was largely aimed at adults during the fall, possibly because of the greater damage they cause to fishing gear and catch. Studies of grey seals have shown that some adults, typically males, may specialize in raiding fishing gear (konigson2013). In contrast, the adult bias was completely absent during the spring hunt. This is not particularly surprising considering that adult seals are molting on ice in large numbers during the spring, and seldom venture out at sea (smith1970; Harkonen1992; born2002; Kelly2010). Younger seals may be much more likely to be foraging in open water during most of the molting period, and thus have more frequent interactions with fishermen (smith1970; lydersen1993; crawford2012; Oksanen2015). In fact, we estimated that adult seals were generally more likely to be hunted than sub-adults during the spring in Finland, where hunting primarily takes place on sea ice. Pups were an exception and were overrepresented in Finnish harvests despite their relatively low presence on sea ice, although some late born pups are often still with their mothers during the Finnish hunting season. Since hunting in Finland is recreational, hunters may specifically seek out inexperienced pups for their meat and fur (kingsley1998).
A population’s response to harvesting is strongly influenced by the age and sex composition of the harvest, with the removal of adult females typically having the greatest impact in long-lived mammals like ringed seals (Kokko1999; heppell2000; Reimer2019). Hence, it is important to ensure that recovered seals are either sampled randomly from the harvests or that any sampling biases are accounted for. In Sweden, hunter-reported data on the sex and body length of each seal, along with records of whether or not each harvested seal was recovered as a sample, allowed us to quantify size- and sex-related sampling biases. Although such biases are unlikely in Finland, similar data could help ensure accurate and unbiased estimates of harvest compositions.
@@
We were not able to estimate the overall magnitude of bycatch mortality due to a lack of data. Such data could be obtained through interviews with fishermen (vanhatalo2014; tubbs2024). However, consistent with findings from many other seal populations, we found that pups were significantly more at risk of getting bycaught in small fisheries than older seals (sipila1990; backlin2011; niemi2013; vanhatalo2014; jounela2019; jounela2024). As pups tend to favor shallow waters for foraging, and are generally curious and inexperienced, it is probable that they frequently engage with coastal fisheries and become caught in fishing gear (Kelly2010; niemi2013; cronin2014). Their relatively small body sizes may also make them more likely to enter smaller fishing gear, and make escaping more difficult once they are caught (cronin2014). The juxtaposition between the substantial juvenile bias in bycatch and the pronounced adult bias in protective hunting by fishermen may reflect the fact that young, inexperienced seals are inclined towards exploration, whereas experienced adults prioritize exploitation (sjoberg2000). On the other hand, larger seals may be more likely to detach from fishing nets, and fishermen often report the difficulty of lifting out larger seals from the water (cronin2014). Older seals may therefore be underrepresented in the samples compared to the actual bycatch.
@@ 5 Conclusions
Amid the dynamic and uncertain conditions brought on by climate change, varying reproductive rates, and the re-introduction of seal hunting, our Bayesian IPM succeeded in providing a comprehensive assessment of the Baltic ringed seal population in the Bothnian Bay – a task that has not been possible for over a decade. In addition to estimating population size and growth, we inferred key demographic parameters including survival probabilities, reproductive rates and the relative vulnerabilities of different demographic groups to hunting and bycatch. Moreover, our model revealed possible changes in ringed seal haul-out behavior in response to changing sea ice patterns. The mechanistic nature of our model additionally enabled near-term predictions of population response to changes in hunting quotas, informing management decisions.
@@ Acknowledgements
The present work is a spin-off from Murat Ersalman’s Master’s thesis (ersalman2024). The work received funding from the Helsinki Institute of Life Science (HiLIFE) (ME), Research Council of Finland (grant 317255) and Jane & Aatos Erkko Foundation (JV). In addition, JV acknowledges funding from the European Union (ERC Consolidator Grant BEFPREDICT, 101087409). The monitoring of ringed seals in the Bothnian Bay was conducted by the Swedish Museum of Natural History and funded by the Swedish Environmental Protection Agency and the Swedish Agency for Marine and Water Management. The Finnish monitoring of seals was funded by the Ministry of Agriculture and Forestry. We thank Yessenia Rojas, Jannikke Räikkönen, Petri Timonen and Charlotta Moraeus for help in age determinations, and Eva Kisdi for support in building the mathematical models. Finally, we are grateful to the two anonymous reviewers for their valuable insights and constructive feedback, which have greatly strengthened this manuscript.
Code and data availability
Code and data are available on Zenodo at: https://doi.org/10.5281/zenodo.14243458
@@
Table 1: Notation for commonly referenced parameters.
Symbol Description
$n_{s,a,t}$ Number of seals of sex $s$ and age $a$ at the start of year $t$
$N_{t}$ Total population size at the start of year $t$
$\rho^{X}_{s,a,t}$ Probability that a seal of sex $s$ and age $a$ transitions to mortality state $X$
$u^{(1)}_{s,a,X,t}$ Number of seals of sex $s$ and age $a$ that transition to mortality state $X$
$u^{(2)}_{s,a,t}$ Number of seals of sex $s$ and age $a$ at the end of year $t$
$H^{i}_{t}$ Harvest totals of country $i$ in year $t$
$\mu_{s,a}$ Mortality rate of seals of sex $s$ and age $a$
$\phi_{s,a}$ Survival probability of seals of sex $s$ and age $a$ in the absence of hunting
$\omega_{s,a,t}$ Haul out probability of seals of sex $s$ and age $a$ in year $t$
$\hat{\omega}_{s,a}$ Haul out probability of seals of sex $s$ and age $a$ when sea ice is abundant
$C_{t}$ Ice extent in the Bothnian Bay in late April
$b_{t}$ Probability that an adult female gives birth at the start of year $t$
@@
Figure 2: Illustration of the Baltic ringed seal annual life-cycle. The density plots at the center show, in polar coordinates, the temporal distribution of hunting and bycatch based on sampled seals. (The illustration was created with the aid of DALL-E 3.)
@@
Figure 4: Estimated LABEL:sub@fig:pop_size population size, LABEL:sub@fig:birth_rate birth rate, LABEL:sub@fig:growth_rate population growth rate and LABEL:sub@fig:age_str age structure for ringed seals in the Bothnian Bay. The rightmost segments show the expected asymptotic values for an unexploited population with and without density dependent effects.
Figure 5: LABEL:sub@fig:survival_probs Probabilities of surviving natural mortality (including bycatch), and the relative probabilities that seals are LABEL:sub@fig:bycatch_bias bycaught, LABEL:sub@fig:sw_bias_spring hunted in Sweden during the spring, LABEL:sub@fig:sw_bias_fall hunted in Sweden during the fall, and LABEL:sub@fig:fi_bias hunted in Finland. The dashed lines in LABEL:sub@fig:bycatch_bias-LABEL:sub@fig:fi_bias show the relative mortality probabilities if all individuals were equally likely to be hunted or bycaught. Calculation of the relative mortality probabilities are presented in Supplement 3.
@@
Figure 7: LABEL:sub@fig:critical_hunting Critical harvest level for which the population growth rate is expected to be zero, as a function of total population size. The critical harvest level is calculated by setting the net reproduction rate equal to one. The red error bar shows the estimated population size and hunting quota (Finland and Sweden combined) in 2022. LABEL:sub@fig:future_preds Posterior predictions for future population growth rates under four different hunting quota scenarios: (1) quotas maintained at their current level of 350 and 375 seals in Sweden and Finland respectively, (2) quotas increased by a total of 30 licenses per year (15 in both Sweden and Finland), (3) quotas increased by a total of 70 licenses per year (35 in both Sweden and Finland) and (4) quotas reduced by a total of 540 in 2024 (270 in both Sweden and Finland) and kept constant thereafter.
Appendix Supplement 1 Model assessment
Figure S1: Posterior predictive check for aerial survey data.
Figure S2: Posterior predictive check for the reproductive sign parameters. Posterior median estimates and central 95% credible intervals for the (population level) proportion of adult females with a LABEL:sub@fig:preg_rate visible fetus, LABEL:sub@fig:scar_rate placental scar, and LABEL:sub@fig:CA_rate corpus albicans are denoted by the black line and grey shading respectively. The observed sample proportions are denoted with dots, whose size reflects the sample size, and the 95% predictive intervals for the sample proportions are shown by the error bars.
@@
Figure S3: Posterior predictive checks for the LABEL:sub@fig:hbag_sw Swedish and LABEL:sub@fig:hbag_fi Finnish ringed seal harvest totals.
@@
Figure S4: Posterior predictive checks for samples from LABEL:sub@fig:h_samples_sw Swedish harvests, LABEL:sub@fig:h_samples_fi Finnish harvests and LABEL:sub@fig:bycatch_samples bycatch. The density estimates are based on the frequency distribution of the number of sampled seals, pooled across all years and demographic classes.
Figure S5: Prior and posterior distributions and overlaps for survival parameters. Dashed lines represent prior distributions.
Figure S6: Prior and posterior distributions and overlaps for haul-out parameters. Dashed lines represent prior distributions.
Figure S7: Prior and posterior distributions and overlaps for reproduction parameters. Dashed lines represent prior distributions.
Figure S8: Prior and posterior distributions and overlaps for hunting and bycatch bias parameters. Dashed lines represent prior distributions. Only parameters related to pups are shown due to space constraints.
Figure S9: Prior and posterior distributions and overlaps for (a) the initial population size and (b) the variance parameter for aerial surveys. Dashed lines represent prior distributions.
Figure S10: Pareto-k statistics for different data types, measuring the relative influence of each data point on the posterior density. (a) The average k-statistic per data point and (b) the sum of k-statistics for each data type, following tinker2024.
Figure S11: Posterior samples of age-specific survival probabilities. Samples are colored to highlight posterior correlations.
Table S1: Posterior overlaps of each model fit with the posterior distribution of the full model. The models are (1) thinned aerial survey data, (2) pre-2012 survey data, (3) post-2012 survey data, (4) thinned reproduction data, (5) post-2007 reproduction data, (6) thinned harvest total data, (7) thinned dead sample data, (8) linearly decreasing reproductive failure rate, (9) left-shifted prior for $\hat{\omega}$, and (10) right-shifted prior for $\hat{\omega}$. Parameter descriptions are given in Supplement 2: Table LABEL:table:priors.
Parameter 1 2 3 4 5 6 7 8 9 10
$n_{1988}$ 0.54 0.41 0.22 0.93 0.62 0.96 0.97 0.98 0.75 0.74
$\phi_{f,5+}$ 0.92 0.94 0.87 0.98 0.85 0.98 0.95 0.99 0.99 0.98
$\phi_{f,0}/\phi_{f,5+}$ 0.81 0.79 0.64 0.97 0.92 0.97 0.97 0.97 0.98 0.98
$\nu_{0}$ 0.98 0.98 0.97 0.98 0.96 0.98 0.98 0.98 0.99 0.98
$\nu_{1+}$ 0.98 0.98 0.96 0.98 0.98 0.98 0.98 0.98 0.99 0.98
$c$ 0.96 0.99 0.97 0.99 0.98 0.99 0.97 0.98 0.98 0.99
$p_{Q}^{\text{fi}}$ 0.75 0.80 0.92 0.98 0.95 0.55 0.96 0.98 0.98 0.96
$p_{Q}^{\text{sw(1)}}$ 0.90 0.51 0.67 0.98 0.92 0.90 0.95 0.98 0.92 0.89
$p_{Q}^{\text{sw(2)}}$ 0.91 0.60 0.72 0.98 0.94 0.50 0.94 0.97 0.91 0.91
$\sigma_{\text{E}_{\text{fi}}}$ 0.93 0.84 0.82 0.97 0.95 0.27 0.95 0.98 0.98 0.99
$\sigma_{\text{E}_{\text{sw(1)}}}$ 0.98 0.98 0.96 0.98 0.97 0.88 0.98 0.98 0.98 0.99
$\sigma_{\text{E}_{\text{sw(2)}}}$ 0.97 0.96 0.96 0.98 0.97 0.50 0.95 0.98 0.98 0.98
$\boldsymbol{\psi}^{\text{bc}}$ 0.88-0.96 0.87-0.95 0.82-0.94 0.98-0.99 0.94-0.98 0.97-0.98 0.78-0.95 0.97-0.99 0.97-0.98 0.97-0.99
$\boldsymbol{\psi}^{\text{sw(1)}}$ 0.88-0.97 0.83-0.97 0.69-0.92 0.97-0.99 0.97-0.99 0.95-0.98 0.56-0.93 0.98-0.99 0.97-0.99 0.98-0.99
$\boldsymbol{\psi}^{\text{sw(2)}}$ 0.88-0.98 0.86-0.98 0.70-0.98 0.97-0.99 0.97-0.98 0.83-0.98 0.66-0.94 0.98-0.99 0.97-0.99 0.97-0.99
$\boldsymbol{\psi}^{\text{fi}}$ 0.77-0.98 0.52-0.87 0.60-0.84 0.97-0.99 0.96-0.98 0.81-0.93 0.71-0.93 0.96-0.98 0.97-0.98 0.95-0.97
$\hat{\omega}$ 0.78 0.70 0.67 0.97 0.94 0.96 0.97 0.98 0.68 0.62
$\delta$ 0.91 0.61 0.84 0.98 0.97 0.90 0.90 0.98 0.97 0.95
$\alpha_{0}/\alpha_{1}$ 0.38 0.22 0.26 0.97 0.97 0.95 0.92 0.97 0.92 0.88
$\alpha_{1}$ 0.50 0.14 0.33 0.98 0.96 0.97 0.96 0.98 0.95 0.94
$f$ 0.57 0.79 0.51 0.97 0.83 0.98 0.96 0.98 0.88 0.87
$d$ 0.58 0.28 0.49 0.98 0.89 0.98 0.98 0.98 0.90 0.93
$b_{\text{max}}$ 0.95 0.90 0.88 0.76 0.85 0.95 0.97 0.80 0.96 0.97
$b_{\text{min}}/b_{\text{max}}$ 0.84 0.94 0.83 0.83 0.54 0.98 0.91 0.99 0.98 0.99
$\beta_{0}$ 0.93 0.90 0.84 0.69 0.40 0.92 0.91 0.96 0.97 0.97
$\beta_{1}$ 0.87 0.92 0.87 0.86 0.44 0.95 0.95 0.96 0.97 0.97
$K$ 0.98 0.98 0.98 0.98 0.98 0.98 0.98 0.98 0.98 0.98
$\theta_{0}$ 0.99 0.96 0.98 0.75 0.97 0.98 0.98 0.77 0.98 0.98
$1/\sqrt{r}$ 0.45 0.77 0.19 0.98 0.85 0.97 0.96 0.98 0.95 0.90
$\pi_{\text{s}}$ 0.95 0.98 0.98 0.77 0.86 0.98 0.99 0.93 0.98 0.98
$\pi_{\text{c}}$ 0.98 0.98 0.99 0.72 0.98 0.97 0.98 0.95 0.98 0.98
$\kappa$ 0.92 0.98 0.97 0.90 0.77 0.98 0.99 0.91 0.98 0.98
Appendix Supplement 2 Prior distributions
@@
We parameterized the age and sex specific median harvest rates, $\hat{E}_{s,a}$, as the product of a global scaling parameter, $\hat{E}$, and a vector of age-specific weights, $\boldsymbol{\psi}$, defined over the unit simplex, so that
@@
Decoupling the magnitude of the harvest rate from its allocation to different demographic groups significantly improves posterior sampling performance. Using Supplement 3: Equation (S7), the scaling parameter $\hat{E}$ can be reparametrized as
@@
Annual variation in harvest rate was modeled as
@@
We chose a large standard deviation of $\sigma=5$ for all components of $\mathbf{g}$ (note that $\sigma$ is on a logarithmic scale in Equation (S6)), which yields a correlation coefficient of $\varrho=0.9887$.
Since the logistic transformation in Equation (S6) requires only 11 degrees of freedom for the 12 demographic groups in our model, we arbitrarily set $g_{12}=0$, and computed the conditional covariance matrix:
@@
We can ensure that the maximum rate of change along the logistic curve in Equation (13) does not exceed the change in pregnancy rate implied by Equation (S14) by satisfying the inequality
@@
Supplement 4: Equation (S4) implies that $e^{-\theta_{0}}$ cannot be less than $b_{\text{max}}$. We therefore gave a Uniform($b_{\text{max}}$, 1) prior to $e^{-\theta_{0}}$. Instead of assigning a prior distribution to $\theta_{1}$, we parametrized our model in terms of the carrying capacity, $K$ (see Supplement 4: Equation (S8)). Estimates of the total number of Baltic ringed seals at the turn of ${}^{\text{th}}$ century range from 50,000 to 450,000 (durant1986; Harding1999; Kokko1999). However, estimates on the lower end of this range are based on very conservative assumptions, including the absence of density dependent regulation. The carrying capacity of the Baltic Sea is therefore highly likely to exceed 200,000 ringed seals (Harding1999). We assumed that, in the absence of hunting pressure, the carrying capacity of the Baltic Sea as a whole is likely to be between 200,000 and 450,000 ringed seals. Based on the distribution of sea ice across the Baltic Sea in late April, about 75% of the Baltic ringed seal population may be present in the Bothnian Bay during the aerial surveys. We therefore gave a Gamma(18, 8e-5) prior to the carrying capacity, $K$, of the Bothnian Bay, which has an expectation of 225,000 (and 95% CI: 133,000 - 340,000).
@@
Survival prob. of female pups $\phi_{f,0}$ $\frac{\phi_{f,0}}{\phi_{f,5+}}\sim U(0,1)$ 0.63 (0.42-0.87)
Deviation between mortality rates of male and female pups $\nu_{0}$ $\nu_{0}\sim N(0,0.1^{2})$ 0.01 (-0.18-0.20)
Deviation between mortality rates of 1+ yo male and females $\nu_{1+}$ $\nu_{1+}\sim N(0,0.1^{2})$ 0.00 (-0.19-0.18)
Shape param. for survival curve $c$ $c\sim U(0,1)$ 0.49 (0.03-0.97)
Scale of FI harvest rate $\hat{E}_{\text{fi}}$ $p_{Q}^{\text{fi}}\sim U(0,1)$ 0.45 (0.26-0.82)
Scale of spring SW harvest rate $\hat{E}_{\text{sw(1)}}$ $p_{Q}^{\text{sw(1)}}\sim U(0,1)$ 2e-3 (6e-4-4e-3)
Scale of fall SW harvest rate $\hat{E}_{\text{sw(2)}}$ $p_{Q}^{\text{sw(2)}}\sim U(0,1)$ 1e-3 (4e-4-4e-3)
SD for FI harvest rate $\sigma_{\text{E}_{\text{fi}}}$ $\sigma_{\text{E}_{\text{fi}}}\sim\text{Cauchy}_{+}(0,0.1)$ 0.48 (0.19-1.29)
SD for spring SW harvest rate $\sigma_{\text{E}_{\text{sw(1)}}}$ $\sigma_{\text{E}_{\text{sw(1)}}}\sim\text{Cauchy}_{+}(0,0.1)$ 1.0 (0.6-2.0)
SD for fall SW harvest rate $\sigma_{\text{E}_{\text{sw(2)}}}$ $\sigma_{\text{E}_{\text{sw(2)}}}\sim\text{Cauchy}_{+}(0,0.1)$ 1.2 (0.6-3.3)
@@
Ratio of pup to adult haul-out $\delta$ $\delta\sim U(0,1)$ 0.58 (0.20-0.94)
Intercept for haul-out model $\alpha_{0}$ $\frac{\alpha_{0}}{\alpha_{1}}\sim\text{Cauchy}(0,5)$ 17.6 (6.2-209.7)
Slope for haul-out model $\alpha_{1}$ $\alpha_{1}\sim\text{Cauchy}(0,1.25)$ 11.0 (4.5-130.7)
Half-saturation const. $f$ $f\sim\text{Cauchy}_{+}(0,37)$ 11.9 (0.9-22.9)
Ratio btw. asymptotes of movement rate into the water $d$ $d\sim U(0,1)$ 0.03 (0.00-0.19)
Max. birth rate $b_{\text{max}}$ $b_{\text{max}}\sim U(0,1)$ 0.77 (0.69-0.87)
Min. birth rate $b_{\text{min}}$ $\frac{b_{\text{min}}}{b_{\text{max}}}\sim U(0,1)$ 0.25 (0.02-0.51)
Intercept for logistic birth rate $\beta_{0}$ $\beta_{0}\sim\text{Cauchy}(0,10)$ 1.7 (-0.5-5.0)
Slope for logistic birth rate $\beta_{1}$ $\beta_{1}\sim\text{t}_{\nu=5}(0,1.8)$ 2.6 (0.9-7.5)
@@
Base pregnancy failure rate $\theta_{0}$ $e^{-\theta_{0}}\sim U(b_{\text{max}},1)$ 0.10 (0.01-0.25)
Dispersion parameter $r$ $\frac{1}{\sqrt{r}}\sim\text{Cauchy}_{+}(0,1)$ 33.0 (17.1-56.6)
Placental scar detection prob. $\pi_{\text{s}}$ $\pi_{\text{s}}\sim U(0,1)$ 0.93 (0.85-0.99)
CA detection prob. $\pi_{\text{c}}$ $\pi_{\text{c}}\sim U(0,1)$ 0.97 (0.91-1.00)
CA prob. for non-pregnant seals $\kappa$ $\kappa\sim U(0,1)$ 0.09 (0.00-0.26)
Sampling prob. for seals $<$100cm $x_{0}$ $x_{0}\sim U(0,1)$ See Supplement 5.
Sampling prob. for seals $>$100cm $x_{1}$ $x_{1}\sim U(0,1)$ See Supplement 5.
Probability that seal is $>$100cm $q_{s,a}$ $q_{s,a}\sim U(0,1)\ \ \forall s,a$ See Supplement 5.
Appendix Supplement 3 Hunting mortality